Sarita bought 2 / 5 metres of ribbon and Lalita 3 /4 metres of ribbon. What is the total length of the ribbon they bought?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Ribbon length bought by Sarita = 2 / 5 metre Ribbon length bought by Lalita = 3 / 4 metre
Total length of the ribbon bought by both of them = 2 / 5 + 3 / 4
Taking LCM 20
= [(2× 4) + (3 × 5)] / 20
= (8 + 15) / 20
= 23 / 20 metre
∴Total length of the ribbon bought by both Sarita and Lalita is 23 / 20 metre

​Thanks